# Atmospheric Density-Compensating Model Predictive Control for Targeted Reentry of Drag-Modulated Spacecraft

### v1

@@ 3.3.3 Measurement Update
where $\mbf{H}=\begin{bmatrix}\mbf{1}_{4\times 4}&\mbf{0}_{4\times 1}\end{bmatrix}$ and $\mbf{V}=\mathbb{E}\{\mbf{v}_{k}\mbf{v}_{k}^{\mathsf{T}}\}$. The matrix $\mbf{H}$ maps the state vector to relative position and velocity. The matrix $\mbf{V}$ is the GPS velocity and position measurement error covariance. The updated state estimate, including the measurement, is then found by
@@ 3.4 Model Predictive Control
In order to guide the spacecraft along the guidance trajectory and to the desired location at the entry interface altitude, the spacecraft state, as estimated by the EKF, is used as feedback within a MPC framework [31]. In this framework, estimated density prediction error is included in the MPC calculation as a predicted disturbance in order to compensate for differences between the density used in generating the guidance trajectory and the density encountered by the spacecraft in flight.
@@
Previous work [31] used a linear time-invariant (LTI) model where the discrete-time dynamics in (49) are computed using $\rho_{\text{nom},0|t}$ and $\Delta\rho_{0|k}$ and assumed constant over the prediction horizon. However, the nominal density, and to a lesser extent, the spacecraft velocity change in time. Therefore, an LTV model of the spacecraft dynamics, where $\mbf{A}_{k}$, $\mbf{B}_{\text{MPC},1,k}$, $\mbf{B}_{\text{MPC},2,k}$ are computed for each time step in the prediction horizon is used to provide a better representation of the relative motion of the spacecraft. Additionally, the density prediciton error appears both in the state space matrices and as a disturbance. If the system is assumed LTI, then $\rho_{\text{nom},0|t}$ and $\Delta\rho_{0|t}$ are used to compute the state space matrices where as $\rho_{\text{nom},k|t}$ and $\Delta\rho_{k|t}$ are used as the disturbance which represents an inconsistency where (10) is computed using values at different points in time. The LTV formulation ensures that the dynamics matrices are computed using values that are consistent in timing with the density prediction error.
Because the spacecraft state is expressed as a position and velocity relative to the guidance trajectory, a feedback controller should drive these states to zero in order to drive the spacecraft to the guidance trajectory. A quadratic cost function of the state is used as a measure of how far the spacecraft is from the guidance during the prediction horizon. The cost function can also penalize the use of control input as was done in previous work [31], where changes in ballistic compared to the nominal are penalized. Similarly, Omar [27] penalizes changes in the ballistic coefficient in order to control when saturation of the drag device occurred. However, the control input is a change in ballistic coefficient from the nominal and no energy is required for a drag device to hold a particular ballistic coefficient. Energy is only required to run the actuator to change the ballistic coefficient. Therefore, because MPC already explicitly accounts for the effect of saturation, it is more useful to penalize changes in the control input between time steps in order to reduce excessive actuation. The cost function that penalizes state error and change in control input is
@@
These bounds are also applicable to the discrete time optimization of (50) if the guidance ballistic coefficient is constant throughout a time step. However, if the guidance ballistic coefficient changes during a time step, a conservative bound is applied to ensure the control limits are not violated. The maximum allowable change in ballistic coefficient is
$\Delta C_{b,\text{max},k|t}=C_{b,\text{max}}-C_{b,g,\text{max},k|t},$ (53)
which is a function of the maximum guidance ballistic coefficient throughout the time step and is given by
$C_{b,g,\text{max},k|t}=\max(C_{b,g}(t))\,\,\text{for all}\,\,t\in[t_{k|t},t_{k+1|t}].$ (54)
Similarly, the conservative lower bound on the change in ballistic coefficient is
$\Delta C_{b,\text{min},k|t}=C_{b,\text{min}}-C_{b,g,\text{min},k|t},$ (55)
where
$C_{b,g,\text{min},k|t}=\min(C_{b,g}(t))\text{for all }t\in[t_{k|t},t_{k+1|t}]$ (56)
is the smallest value of the guidance ballistic coefficient throughout the $k^{th}$ time step. The control inputs can then be constrained by
@@
The optimization consisting of the minimization of (50) subject to the constraints of the control input given by (57) is formulated as a quadratic program, which is convex. This allows the problem to be solved quickly and efficiently using an off-the-shelf solver. The built-in MATLAB solver, quadprog.m, is used to solve the optimization problem in this work, however other solvers such as OSQP [43] or custom code generated by CVXgen [44] are suitable for real-time use in an embedded application such as onboard a spacecraft.
@@ 4 Simulations
To evaluate the performance of the proposed estimation and control framework under realistic flight conditions, a Monte-Carlo campaign consisting of 1000 runs is conducted. Each run generates a guidance trajectory beginning in a random orbit at approximately $375$ km in altitude and ending at a random latitude and longitude at the von Karman altitude of $100$ km. The estimation and control framework is then simulated in a custom simulation environment beginning at a dispersed initial position and velocity and in the presence of atmospheric density errors, gravitational modeling errors and measurement noise. The simulation continues until the spacecraft reaches $100$ km in altitude and the distance by which the spacecraft missed the desired reentry location is recorded. Should the tracking spacecraft drift away from the guidance trajectory by more than $100$ km at any point in time, the simulation is terminated and the tracking is considered to have failed. The simulation environment, along with the EKF and MPC, are depicted in Fig. 4. For comparative purposes, the estimation and control framework is simulated twice for each trajectory. One simulation makes use of the estimated density error within MPC, while the other does not.
Figure 4: Block diagram depicting the simulation of estimation and control algorithms and the spacecraft dynamics. The upper box depicts the computations that would occur onboard the spacecraft where as the lower box depicts the simulated spacecraft environment.
@@ 4.1 Environmental Modeling
A custom simulation environment is developed in MATLAB leveraging the high-fidelity models included in the Aerospace Toolbox [45]. These include models of Earth’s gravity and atmosphere, which are used to calculate the different forces that influence the motion of the spacecraft. The dynamics of the spacecraft are propagated using the variable-step ode113.m solver. This custom simulation environment has been validated against multiple simulation test cases provided by the NASA Engineering and Safety Center [46], as shown in [47, Section 5.2.2].
4.1.1 Geodesy
The Earth is modeled as an oblate spheroid using the WGS84 [48] ellipsoid. The altitude of the spacecraft is then calculated as the height of the spacecraft above this ellipsoid. The Earth is assumed to rotate about the $\underrightarrow{ECI}^{3}$ direction resulting in an angular velocity of the $ECEF$ frame relative to the $ECI$ frame of ${\boldsymbol{\omega}}^{ECEF-ECI}_{ECI}=\begin{bmatrix}0&0&7.292115\times 10^{-5}\end{bmatrix}^{\mathsf{T}}$ rad/s. The orientation of the Earth at the initial time is obtained from the International Astronomical Union - 2000/2005 reference system as implemented in the MATLAB function dcmeci2ecef.m [45]. The direction cosine matrix obtained from this function is converted to an Euler angle representation. The angle corresponding to rotation about the $\underrightarrow{ECI}^{3}$ axis is the only angle considered. The other two angles are assumed to be zero, which introduces the assumption that the $\underrightarrow{ECI}^{3}$ and $\underrightarrow{ECEF}^{3}$ axes are aligned for the duration of the simulation.
4.1.2 Gravity
The gravitational acceleration acting on the spacecraft is modeled using the EGM2008 [49] model as implemented in the MATLAB function gravitysphericalharmonic.m [45]. This model uses a spherical harmonic representation of the gravitational potential of the Earth. The fidelity of this model is adjustable by considering a different number of harmonic terms. In order to test the presented estimation and control framework in the presence of gravity modeling errors, the guidance trajectory is generated using the gravity model through a degree and order of 15 where as the tracking simulations are performed with gravity modeled through a degree and order of 40. The additional gravitational terms used in the tracking introduce a dispersion to the nominal gravity used in generating the guidance trajectory.
4.1.3 Atmospheric Drag
@@
$\mbf{v}^{\text{atmo},p/ECI}_{ECI}={\boldsymbol{\omega}}^{ECEF-ECI^{\times}}_{ECI}\mbf{r}^{sp}_{ECI}.$ (58)
@@
The magnitude of this relative velocity is used in computing the magnitude of the drag acceleration where as the direction of the relative velocity is the direction in which the drag acceleration acts.
The density of the atmosphere is obtained from the NRLMSISE00 [34] atmosphere model as implemented in atmosnrlmsise00.m [45]. In the tracking simulations, historical observations [50] of the $F10.7$ solar flux and $A_{p}$ geomagnetic index are used when evaluating the atmosphere model in order to provide realistic variations in the density. In order to introduce errors in the density prediction between the guidance trajectory and the tracking simulation, imperfect knowledge of the space weather environment are used when generating the guidance. During the guidance generation, the 81-day centered average of the $F10.7$ solar flux is used in place of the daily observed values. In place of the $A_{p}$ geomagnetic index for the current time, the average of eight 3-hour $A_{p}$ observations from 36 to 57 hours before the current time is used. These averages are used to represent the imperfect knowledge of future space weather indices that would be available in the form of forecasts.
The discrepancy in the space weather indices between the guidance and the tracking is used to produce density prediction errors representative of those that result from the errors in space weather forecasting, but there are other sources of uncertainty in the drag force, such as uncertainty in the atmosphere model itself, as well as uncertainty in the area and drag coefficient of the drag device. For the tracking simulations, a constant multiplier $c_{\text{drag}}$ is added to the drag acceleration to represent additional sources of drag error. This multiplier is randomly sampled between individual simulations.
@@ 4.2 Monte Carlo Parameters
For each individual simulation, the initial condition of the guidance trajectory is generated by sampling the orbital elements of the initial orbit at a random epoch obtained from uniform distributions. All of these orbits begin at an altitude of approximately $375$ km by choosing an initial semi-major axis of $6853$ km, which was found to be sufficiently high to provide enough longitude controllability to target any point below the inclination of the spacecraft orbit, regardless of the state of the atmosphere. The desired latitude and longitude at entry interface are also sampled and the entry interface altitude is chosen to be $100$ km. The sampled parameters used in generating the guidance are listed in Table 1.
@@
After the guidance trajectory is generated, it is then tracked by the spacecraft. The initial radial and in-track position of the spacecraft at the initial time is perturbed by adding a dispersion to the initial semi-major axis and true anomaly of the initial orbit of the tracking spacecraft corresponding to dispersions up to approximately $100$ m in the radial direction and $1$ km in the in-track direction. The density multiplier, $c_{\text{drag}}$, is sampled form the interval ranging from $0.75$ to $1.25$. The position and velocity of the spacecraft relative to the guidance are corrupted by noise at the beginning of each time step according to (43) to produce synthetic GPS measurements that are used for the measurement update in the EKF. The noise is zero-mean and normally distributed according to the parameters in Table 2
@@ 4.3 Algorithm Parameters
Both the EKF and MPC are run with a sample time of $\Delta T\approx 600$ s which corresponds to a new state estimate and control input being computed and implemented approximately every 10 minutes or 9 times per orbit. The exact sample time is obtained by dividing the duration of the guidance trajectory by the desired sample time to obtain the number of time steps. This number is rounded to the nearest integer to compute the actual sample time. The EKF is run using the parameters in Table 2 while the MPC uses the parameters listed in Table 3.
@@ 4.4 Results
A guidance trajectory was successfully found for all 1000 cases. The desired final latitude, longitude and altitude for each case are used to obtain the desired final position of the spacecraft in the $ECEF$ frame. The targeting error of each guidance is then computed as the norm of the difference between desired final position of the spacecraft and final position of the guidance trajectory, both in the $ECEF$ frame. A histogram of the targeting errors of each guidance trajectory is plotted in Fig. 5a. The mean targeting error is $6.31$ km and $97.7\%$ of cases have targeting error below the tolerance of $10$ km while $99.5\%$ of cases have targeting errors below $100$ km. The maximum targeting error is $262$ km. The ballistic coefficients that produce each guidance trajectory are shown in Fig. 5b.
@@
Figure 5: Monte Carlo results of guidance trajectory generation, including (a) final position errors of the guidance trajectories and (b) nominal ballistic coefficients for the guidance trajectories (blue) and the spacecraft ballistic coefficient bounds (dashed black).
@@
The guidance trajectories are tracked down to the desired entry interface altitude of $100$ km using the proposed estimation and control framework. If the spacecraft reaches the final altitude before the end of the guidance trajectory, then the simulation is terminated early when the spacecraft is at an altitude of $100$ km. Should the spacecraft be above the desired altitude when the guidance trajectory ends, the simulation continues until the spacecraft reaches the desired final altitude. The final tracking error at entry interface is then computed as the difference between the final position of the guidance and the final position of the spacecraft at the entry interface altitude, both in the $ECEF$ frame. Computing the final tracking error in this way is more indicative of the reentry accuracy compared to looking at the radial and in-track errors at the final time of the guidance trajectory. This is because tracking error in the radial and in-track directions at the final time of the guidance trajectory contribute differently to the reentry accuracy. Errors in the in-track direction at the final time of the guidance produce errors of similar magnitude throughout the trajectory after entry interface, whereas errors in the radial direction are magnified. By computing errors at the entry interface altitude instead of at the final time of the guidance, the relative contribution of in-track and radial errors at the final time on the entry accuracy is captured.
@@
Of the $1000$ simulations, 984 trajectories were successfully tracked down to entry interface in both the case where $\Delta\rho$ estimates were and were not used in the MPC controller, leading to a success rate of $98.4\%$. In the other $16$ cases, the tracking error exceeded the $100$ km threshold at some point in time and the simulations were terminated early. The final tracking errors at entry interface are shown in Fig. 6. Without including the estimates of $\Delta\rho$ in the control, the mean tracking error was $244$ km with $23.0\%$ of cases falling below 100 km. Including estimates of $\Delta\rho$ results in a mean tracking error of $12.1$ km, where $99.7\%$ cases fall below 100 km.
@@
Monte Carlo simulations with the estimates of $\Delta\rho$ are repeated for three different values of $R_{c}$ in the MPC controller to observe the effect of tuning on the control effort and final tracking error. Only the first $500$ randomly-generated trajectories are considered and the gravity model is set to the same fidelity used during the guidance generation to decrease run time. To obtain a sense of how much the drag device is used, the sum of all of the $\Delta C_{b}$ for a particular trajectory is divided by the duration of the corresponding trajectory. Fig. 7a shows the actuator use per unit time for all of the trajectories. Fig. 7b shows a histogram of the final tracking errors for the three different tunings.
@@
Figure 6: Tracking errors at the entry interface altitude with (red) and without (blue) estimated $\Delta\rho$ information in the control.
@@
Figure 7: Results with different tunings of the MPC controller, including (a) actuator use per unit time and (b) tracking errors at the entry interface altitude.
The true value of $c_{\Delta\rho}$ for one particular case from the Monte Carlo is shown in Fig. 8 along with the value estimated by the EKF. The control inputs produced by MPC in both the case where the estimated density information is not used and the case where it is used are shown in Fig. 9. The tracking errors produced in these two cases are shown in Fig. 10.
@@
Figure 9: Ballistic coefficient commanded by MPC without (blue) and with (red) estimated density information included in the control computation.
@@
5 Discussion
While the guidance generation algorithm is not a contribution of this work, the simplified implementation of the work by Omar successfully produced a guidance trajectory in all cases with performance comparable to the original implementation [40]. The vast majority of the targeting errors were small enough to safely target reentry away from populated areas, as shown in Fig. 5a. Additionally, our implementation is able to select trajectories corresponding to nominal ballistic coefficients that preserve more control authority to decrease the ballistic coefficient, as shown in Fig. 5b compared to Fig. 3. While there are some outlier trajectories with large targeting errors, this can likely be improved using methods present in Omar’s more sophisticated implementations, such as drag-work enforcement, back stepping and using a shrinking horizon. However, our simplified implementation was sufficient for the purpose of providing trajectories on which to test the proposed estimation and control methods.
The combined MPC and EKF approach was able to guide the spacecraft accurately along the guidance trajectory and to the desired entry interface location in almost every case. Of the cases that were terminated early, several coincided with periods of volatile atmospheric conditions associated with solar storms and flares such as those in 2005. Under such extreme solar conditions, the atmospheric density changes to an extent that the drag control device inevitably becomes saturated and the guidance trajectory is not trackable. The tracking errors for the successful cases, as shown in Fig. 6, were sufficiently small for the spacecraft to safely target a reentry over the ocean with a success rate of $98.4\%$. This demonstrates the capability of the proposed targeted reentry control method for the deorbit of small satellites with mission requirements of safely splashing down in the ocean, such as the HyCUBE concept [20, 21].
As shown in Fig. 6, the inclusion of the estimated density error from the EKF drastically improves the tracking performance of the MPC controller. Looking at the single case in Fig. 8, while not capturing the oscillations that occur with the same period of the spacecraft orbit, the EKF is able to estimate an average value of the normalized density prediction error. This estimated knowledge of the atmosphere is able to remove a significant amount of the tracking error that builds up with out this knowledge, as shown in Fig. 10. This, however, requires the ability to adjust the ballistic coefficient of the spacecraft very precisely. As shown in Fig. 9, the commanded ballistic coefficients are only very slightly different than those commanded without the density error estimate.
Increasing $R_{c}$ to increase the cost of changing the control input reduces the total change in the ballistic coefficient throughout the trajectories, as can be seen in Fig. 7a. However, the reduced actuation results in larger tracking errors when the spacecraft reaches entry interface, as evident in Fig. 7b. This demonstrates that the controller is able to be adjusted to give the desired balance of tracking performance and actuator use.

### v2

@@ 3.3.3 Measurement Update
where $\mbf{H}=\begin{bmatrix}\mbf{1}_{4\times 4}&\mbf{0}_{4\times 1}\end{bmatrix}$ maps the state vector to relative position and velocity, while $\mbf{V}=\mathbb{E}\{\mbf{v}_{k}\mbf{v}_{k}^{\mathsf{T}}\}$ is the GPS velocity and position measurement error covariance. The updated state estimate, including the measurement, is then found by
@@ 3.4 Model Predictive Control
In order to guide the spacecraft along the guidance trajectory and to the desired location at the entry interface altitude, the spacecraft state, as estimated by the EKF, is used as feedback within a MPC framework [33]. In this framework, the estimated density prediction error is included in the MPC calculation as a predicted disturbance in order to compensate for differences between the density used in generating the guidance trajectory and the density encountered by the spacecraft in flight.
@@
Previous work [33] used a linear time-invariant (LTI) model where the discrete-time dynamics in (39) are computed using $\rho_{\text{nom},0|t}$ and $\Delta\rho_{0|k}$ and assumed constant over the prediction horizon. Because the nominal density and, to a lesser extent, the spacecraft velocity are time-varying, an LTV model of the spacecraft dynamics is used, where $\mbf{A}_{k}$, $\mbf{B}_{\text{MPC},1,k}$, $\mbf{B}_{\text{MPC},2,k}$ are computed for each time step in the prediction horizon. Additionally, the density prediction error appears both in the state space matrices and as a disturbance. If the system were assumed to be LTI, then $\rho_{\text{nom},0|t}$ and $\Delta\rho_{0|t}$ would be used to compute the state-space matrices. This would be inconsistent with the use of $\rho_{\text{nom},k|t}$ and $\Delta\rho_{k|t}$, which are time-varying nominal density and density prediction error terms. The LTV formulation ensures that the dynamics matrices are computed using values that are consistent in timing with the density prediction error.
When choosing a cost function for MPC to minimize, it is important to consider the practical needs of the mission. Because the spacecraft state is expressed as a position and velocity relative to the guidance trajectory, a feedback controller should drive these states to zero in order to drive the spacecraft to the guidance trajectory. This motivates the inclusion of a state cost that penalizes how far the spacecraft is from the guidance during the prediction horizon. The cost function should also penalize the use of control input as was done in previous work [33], where changes in ballistic coefficient compared to the nominal were penalized. Similarly, Omar [29] penalized changes in the ballistic coefficient in order to control when saturation of the drag device occurred. However, the control input is a change in ballistic coefficient from the nominal and no energy is required for a drag device to hold a particular ballistic coefficient. Energy is only required to run the actuator to change the ballistic coefficient. Therefore, because MPC already explicitly accounts for the effect of saturation, it is more useful to penalize changes in the control input between time steps in order to reduce excessive actuation.
Motivated by the prior work of [29, 33] and the desire to penalize only a change in control input between time steps, the proposed MPC cost function is
@@
These bounds are also applicable to the discrete time optimization of (40) if the guidance ballistic coefficient is constant throughout a time step. However, if the guidance ballistic coefficient changes during a time step, a conservative bound must be used to ensure the control limits are not violated. The control inputs can then be constrained by
@@
The optimization consisting of the minimization of (40) subject to the constraints of the control input given by (43) is formulated as a quadratic program, which is convex and can be solved quickly and efficiently. For convenience, the built-in MATLAB quadratic program solver quadprog is used to solve the optimization problem in this work, however other solvers such as OSQP [54] or custom code generated by CVXgen [55] are suitable for real-time use in an embedded application such as onboard a spacecraft.
@@ 4 Simulations
To evaluate the performance of the proposed estimation and control framework under realistic flight conditions, a Monte-Carlo campaign consisting of 1000 runs is conducted. Each run generates a guidance trajectory beginning in a random near-circular orbit at approximately $375$ km in altitude and ending at a random latitude and longitude at the von Karman altitude of $100$ km. Although the altitude of entry interface is not precisely defined in the literature [56], $100$ km is chosen in this work to align with prior studies on drag-based deorbit trajectories [24, 57, 58]. The estimation and control framework is then simulated in a custom simulation environment beginning at a dispersed initial position and velocity and in the presence of atmospheric density errors, gravitational modeling errors and measurement noise. The simulation continues until the spacecraft reaches $100$ km in altitude and the distance by which the spacecraft missed the desired reentry location is recorded. Should the tracking spacecraft drift away from the guidance trajectory by more than $100$ km at any point in time, the simulation is terminated and the tracking is considered to have failed. The simulation environment, along with the EKF and MPC, are depicted in Fig. 4. For comparative purposes, the estimation and control framework is simulated twice for each trajectory. One simulation makes use of the estimated density error within MPC, while the other does not.
Figure 4: Block diagram depicting the simulation of estimation and control algorithms and the spacecraft dynamics.
@@ 4.1 Environmental Modeling
A custom simulation environment is developed in MATLAB leveraging the high-fidelity models included in the Aerospace Toolbox [59]. These include models of Earth’s gravity and atmosphere, which are used to calculate the different forces that influence the motion of the spacecraft. The dynamics of the spacecraft are propagated using the variable-step ode113 solver. This custom simulation environment has been validated against multiple simulation test cases provided by the NASA Engineering and Safety Center [60].
4.1.1 Nominal Models
The Earth is modeled as an oblate spheroid using the WGS84 [61] ellipsoid. The altitude of the spacecraft is then calculated as the height of the spacecraft above this ellipsoid. The Earth is assumed to rotate about the $\underrightarrow{ECI}^{3}$ direction resulting in an angular velocity of the $ECEF$ frame relative to the $ECI$ frame of ${\boldsymbol{\omega}}^{ECEF-ECI}_{ECI}=\begin{bmatrix}0&0&7.292115\times 10^{-5}\end{bmatrix}^{\mathsf{T}}$ rad/s. The orientation of the Earth at the initial time is obtained from the International Astronomical Union - 2000/2005 reference system as implemented in the MATLAB function dcmeci2ecef [59] and the nutation and precession is neglected.
The gravitational acceleration acting on the spacecraft is modeled using the EGM2008 [62] model as implemented in the MATLAB function gravitysphericalharmonic [59]. This model is used with degree and order 15 to generate nominal values of the gravitational acceleration.
@@ 4.1.1 Nominal Models
$\mbf{v}^{\text{atmo},p/ECI}_{ECI}=\left({\boldsymbol{\omega}}^{ECEF-ECI}_{ECI}\right)^{\times}\mbf{r}^{sp}_{ECI}.$ (44)
@@
The magnitude of this relative velocity is used in computing the magnitude of the drag acceleration, whereas the direction of the relative velocity is the direction in which the drag acceleration acts.
The density of the atmosphere is obtained from the NRLMSISE00 [40] atmosphere model as implemented in atmosnrlmsise00 [59].To generate nominal values of the density, the 81-day centered average of the $F10.7$ solar flux is used in place of the daily observed values. In place of the $A_{p}$ geomagnetic index for the current time, the average of eight 3-hour $A_{p}$ observations from 36 to 57 hours before the current time is used. These averages are used to represent the imperfect knowledge of future space weather indices that would be available in the form of forecasts when predicting a nominal density profile.
4.1.2 Dispersed Models
When generating guidance trajecories, the nominal environmental models are used. However, in order to evaluate the performance of the estimation and control methods in the presence of modeling errors and uncertainties, dispersed environmental models are used in the tracking simulations.
In order to represent modeling error and uncertainty in the gravitational acceleration, the EGM2008 [62] model is dispersed by using a higher degree and order of 40 compared to the nominal gravity model which produces differences in the effect of gravity between the guidance trajectory and the tracking simulation.
In the tracking simulations, real historical observations [63] of the $F10.7$ solar flux and $A_{p}$ geomagnetic index are used when evaluating the atmosphere model in order to provide realistic variations in the density. The discrepancy in the space weather indices between the guidance trajectory and the tracking simulation is used to produce density prediction errors representative of those that result from the errors in space weather forecasting.
There are other sources of uncertainty in the drag force, such as error in the atmosphere model itself, as well as uncertainty in the area and drag coefficient of the drag device. For the tracking simulations, a constant multiplier $c_{\text{drag}}$ is added to the drag acceleration to represent these additional sources of drag error. This multiplier is randomly sampled between individual simulations and held constant for the duration of each trajectory.
@@ 4.2 Monte Carlo Parameters
The sampled parameters used in generating the guidance trajectory are listed in Table 1. For each individual simulation, the initial condition of the guidance trajectory is generated by sampling the orbital elements of the initial orbit at a random epoch obtained from uniform distributions. All of these orbits begin at an altitude of approximately $375$ km by choosing an initial semi-major axis of $6853$ km, which was found to be sufficiently high to provide enough longitude controllability to target any point below the inclination of the spacecraft orbit, regardless of the state of the atmosphere. The desired latitude and longitude at entry interface are also sampled and the entry interface altitude is chosen to be $100$ km.
@@
After the guidance trajectory is generated, it is then tracked by the spacecraft. The initial radial and in-track position of the spacecraft at the initial time is perturbed by adding a dispersion to the initial semi-major axis and true anomaly of the initial orbit of the tracking spacecraft corresponding to dispersions up to approximately $100$ m in the radial direction and $1$ km in the in-track direction. The density multiplier, $c_{\text{drag}}$, is sampled uniformly from the interval ranging from $0.75$ to $1.25$. The position and velocity of the spacecraft relative to the guidance are corrupted by noise at the beginning of each time step according to (33) to produce synthetic GPS measurements that are used for the measurement update in the EKF. The noise is zero-mean and normally distributed according to the parameters in Table 2 which are chosen to be representative of GPS measurement errors [29, 65].
@@ 4.3 Algorithm Parameters
Both the EKF and MPC are run with a sample time of $\Delta T\approx 600$ s which corresponds to a new state estimate and control input being computed and implemented approximately every 10 minutes or 9 times per orbit. The exact sample time is obtained by dividing the duration of the guidance trajectory by the desired sample time to obtain the number of time steps. This number is rounded to the nearest integer to compute the actual sample time.
The EKF is run using the parameters in Table 2. The matrix $\mbf{V}$ is chosen to match the noise characteristics used in generating the synthetic measurements while the process noise $\mbf{W}$ is treated as a tuning parameter and is selected to provide the desired level of performance in estimating the density error.
The MPC uses the parameters listed in Table 3. The majority of these parameters are controller design parameters that are tuned to give the desired level of tracking performance. The control bounds, $C_{b,\text{min}}$ and $C_{b,\text{max}}$, are physical design parameters that depend on the drag control device being used. The values used here are selected to be representative of the range of ballistic coefficients that are achievable using a device similar to the ExoBrake [4].
@@ 4.4 Results
A guidance trajectory was successfully found for all 1000 cases. The desired final latitude, longitude and altitude for each case are used to obtain the desired final position of the spacecraft in the $ECEF$ frame. The targeting error of each guidance trajectory is then computed as the norm of the difference between desired final position of the spacecraft and final position of the guidance trajectory, both in the $ECEF$ frame. A histogram and cumulative distribution function of the targeting errors of each guidance trajectory are plotted in Fig. 5. The mean targeting error is $6.31$ km and $97.7\%$ of cases have targeting error below the tolerance of $10$ km while $99.5\%$ of cases have targeting errors below $100$ km. The maximum targeting error is $262$ km. The ballistic coefficients that produce each guidance trajectory are shown in Fig. 6. While the guidance generation algorithm is not a contribution of this work, the simplified implementation of the work by Omar successfully produced a guidance trajectory in all cases with performance comparable to the original implementation [48]. The vast majority of the targeting errors were small enough to safely target reentry away from populated areas, as shown in Fig. 5. Additionally, our implementation is able to select trajectories corresponding to nominal ballistic coefficients that preserve more control authority to decrease the ballistic coefficient, as shown in Fig. 6 compared to Fig. 3. While there are some outlier trajectories with large targeting errors, this can likely be improved using methods present in Omar’s more sophisticated implementations, such as drag-work enforcement, back stepping and using a shrinking horizon. However, our simplified implementation was sufficient for the purpose of providing trajectories on which to test the proposed estimation and control methods.
Figure 5: Monte Carlo results of the final guidance position errors as (a) a histogram and (b) a cumulative distribution function zoomed in on the errors below $25$ km.
@@
Figure 6: Nominal ballistic coefficients for the Monte Carlo guidance trajectories (blue) and the spacecraft ballistic coefficient bounds (dashed black).
@@
The guidance trajectories are tracked down to the desired entry interface altitude of $100$ km using the proposed estimation and control framework. If the spacecraft reaches the final altitude before the end of the guidance trajectory, then the simulation is terminated early when the spacecraft is at an altitude of $100$ km. Should the spacecraft be above the desired altitude when the guidance trajectory ends, the simulation continues until the spacecraft reaches the desired final altitude. In this case, the final tracking error at entry interface is then computed based on the final position of the guidance and the final position of the spacecraft at the entry interface altitude. Computing the final tracking error in this way is more indicative of the reentry accuracy compared to looking at the radial and in-track errors at the final time of the guidance trajectory. This is because tracking error in the radial and in-track directions at the final time of the guidance trajectory contribute differently to the reentry accuracy. Errors in the in-track direction at the final time of the guidance produce errors of similar magnitude throughout the trajectory after entry interface, whereas errors in the radial direction are magnified. By computing errors at the entry interface altitude instead of at the final time of the guidance, the relative contribution of in-track and radial errors at the final time on the entry accuracy is captured.
@@
Of the $1000$ simulations, 984 trajectories were successfully tracked down to entry interface in both the case where $\Delta\rho$ estimates were and were not used in the MPC controller, leading to a success rate of $98.4\%$. In the other $16$ cases, the tracking error exceeded the $100$ km threshold at some point in time and the simulations were terminated early. The final tracking errors at entry interface are shown in Fig. 7. Without including the estimates of $\Delta\rho$ in the control, the mean tracking error was $244$ km with $23.0\%$ of cases falling below 100 km. Including estimates of $\Delta\rho$ results in a mean tracking error of $12.1$ km, where $99.7\%$ cases fall below 100 km. The combined MPC and EKF approach was able to guide the spacecraft accurately along the guidance trajectory and to the desired entry interface location in almost every case. Of the cases that were terminated early, several coincided with periods of volatile atmospheric conditions associated with solar storms and flares such as those in 2005. Under such extreme solar conditions, the atmospheric density changes to an extent that the drag control device inevitably becomes saturated and the guidance trajectory is not trackable. The tracking errors for the successful cases, as shown in Fig. 7, were sufficiently small for the spacecraft to safely target a reentry over the ocean with a success rate of $98.4\%$. This demonstrates the capability of the proposed targeted reentry control method for the deorbit of small satellites with mission requirements of safely splashing down in the ocean, such as the HyCUBE concept [22, 23].
@@
Monte Carlo simulations with the estimates of $\Delta\rho$ are repeated for three different values of $R_{c}$ in the MPC controller to observe the effect of tuning on the control effort and final tracking error. Only the first $500$ randomly-generated trajectories are considered and the gravity model is set to the same fidelity used during the guidance generation to decrease run time. To obtain a sense of how much the drag device is used, the sum of all of the $\Delta C_{b}$ for a particular trajectory is divided by the duration of the corresponding trajectory. Fig. 8 shows the actuator use per unit time for all of the trajectories. Fig. 9 shows a histogram and cumulative distribution function of the final tracking errors for the three different tunings. Increasing $R_{c}$ to increase the cost of changing the control input reduces the total change in the ballistic coefficient throughout the trajectories, as can be seen in Fig. 8. However, the reduced actuation results in larger tracking errors when the spacecraft reaches entry interface, as evident in Fig. 9. This demonstrates that the controller is able to be adjusted to give the desired balance of tracking performance and actuator use.
@@
Figure 7: Entry interface tracking errors with (red) and without (blue) estimated $\Delta\rho$ information in the control as (a) a histogram and (b) a cumulative distribution function.
@@
Figure 8: Actuator use per unit time with different tunings of the MPC controller.
Figure 9: Entry interface tracking errors with different tunings of the MPC controller as (a) a histogram and (b) a cumulative distribution function.
The true value of $c_{\Delta\rho}$ for one particular case from the Monte Carlo is shown in Fig. 10 along with the value estimated by the EKF. The control inputs produced by MPC in both the case where the estimated density information is not used and the case where it is used are shown in Fig. 11a. The tracking errors produced in these two cases are shown in Fig. 12. As shown in Fig. 7, the inclusion of the estimated density error from the EKF drastically improves the tracking performance of the MPC controller. The single case in Fig. 10 demonstrates that the EKF is able to estimate an average value of the normalized density prediction error, although it does not capture the oscillations that occur with the same period of the spacecraft orbit. This estimated knowledge of the atmosphere is able to remove a significant amount of the tracking error that builds up with out this knowledge, as shown in Fig. 12. This, however, requires the ability to adjust the ballistic coefficient of the spacecraft very precisely. As shown in Fig. 11a, the commanded ballistic coefficients are only very slightly different than those commanded without the density error estimate. The difference between these ballistic coefficients normalized by the ballistic coefficient without estimated knowledge of the density error is included in Fig. 11b.
@@
Figure 11: (a) Ballistic coefficient commanded by MPC without (blue) and with (red) estimated density information and (b) the normalized difference between them.